# Power to the Clients: Federated Learning in a Dictatorship Setting

### v1

@@ Abstract
1 Introduction
Federated learning (FL) (McMahan et al., 2017) is a distributed learning paradigm in which model training is performed collaboratively by a set of clients. In centralized FL, a global server broadcasts the current model to all clients, each of which updates the model using its local dataset and sends back the resulting gradients to the server. The server then aggregates these gradients to update the global model. This approach accelerates training by distributing computation across multiple machines, while also enhancing data privacy since clients share only gradients, not their raw data. FL is especially well-suited for privacy-sensitive applications, such as training on confidential medical records across hospitals.
Despite its advantages, FL remains vulnerable to malicious behavior by the participating clients. Byzantine clients are adversarial participants that disrupt the training process by sending arbitrary or manipulated updates to the central server (Lamport et al., 2019)(Blanchard et al., 2017). The presence of such adversaries can significantly degrade model performance, making Byzantine robustness a critical area of study (Li et al., 2019; Wu et al., 2020; Shejwalkar & Houmansadr, 2021; Guerraoui et al., 2018; Xie et al., 2018; Xie, 2022). Moreover, several studies have demonstrated the possibility of backdoor attacks in FL via collusion attacks where multiple malicious clients coordinate their actions to inject hidden triggers into the global model in FL (Liu et al., 2024; Ranjan et al., 2022; Xiao et al., 2022; Bagdasaryan et al., 2020). These clients may exchange information and strategically craft updates that steer the aggregated model toward a compromised state.
@@ 1 Introduction
To demonstrate the feasibility of this threat, we develop a series of algorithms that enable malicious clients to achieve their goals within the defined constraints. Our theoretical findings are further supported by empirical results, which validate the effectiveness of the proposed attack strategies. Beyond isolated attacks, we also investigate complex and previously underexamined dynamics that arise among malicious clients themselves. For example, we examine scenarios in which all participants in the system act as dictators, as well as cases where collaborative dictator clients can betray one another within their own partnership. These scenarios reveal internal conflicts among adversaries and broaden the understanding of multi-agent adversarial behavior in FL.
2 Related Work
The distributed nature of FL, combined with the server’s limited visibility into local training processes, makes it vulnerable to various security threats posed by malicious or compromised clients (Zhang et al., 2023b). In this section, we review existing literature across three major category of attacks—Byzantine attacks, backdoor attacks, and collusion attacks.
@@ 2 Related Work
Byzantine attacks pose a fundamental threat in distributed systems including FL, where a subset of clients, known as Byzantine clients, arbitrarily deviate from the prescribed protocol by submitting malicious or anomalous updates to the central server (Lamport et al., 2019). The goals of such attacks typically include degrading the global model’s performance or preventing convergence (Blanchard et al., 2017). Attack strategies vary in complexity, ranging from simple approaches such as random noise injection or submitting zero gradients to more sophisticated methods like sign-flipping (Samy & Girdzijauskas, 2023; Shen et al., 2025). Advanced attacks are often crafted to evade specific defenses, making them challenging to detect and mitigate (Shejwalkar & Houmansadr, 2021; Baruch et al., 2019).
@@
Backdoor attacks (also known as Trojan attacks) are a more insidious threat in FL where attackers aim to embed hidden malicious behavior into the global model (Gu et al., 2017; Li et al., 2022). An attacker, typically controlling one or more clients, manipulates their local dataset or model updates to create a ”backdoor trigger”—a specific pattern or feature (e.g., a small patch in an image, a specific phrase in text). The compromised global model performs normally on clean inputs but exhibits attacker-chosen behavior, such as misclassification, when the trigger is present. These attacks can be implemented through various strategies, including data poisoning, where labels are manipulated for samples containing the trigger, and model poisoning, where malicious updates are directly crafted to influence model behavior (Bagdasaryan et al., 2020; Xie et al., 2020). Triggers may be static and predefined (Bagdasaryan et al., 2020) or dynamically generated using optimization techniques to make them more subtle and difficult to detect (Zhang et al., 2023a). Comprehensive surveys on backdoor attacks and defenses in FL can be found in Nguyen et al. (2024).
@@
Collusion attacks occur when multiple malicious clients coordinate their actions to enhance the effectiveness of the attacks or bypass defenses designed for independent attackers. Colluding attackers can amplify the impact of Byzantine or backdoor attacks. For example, multiple Byzantine clients might coordinate their updates to overwhelm Byzantine-resilient aggregation rules that assume the number of attackers are limited (Xie et al., 2020). Similarly, colluding clients can implement distributed backdoor attacks, where each attacker contributes a part of the malicious update, making individual contributions appear benign while collectively embedding a backdoor into the global model (Lyu et al., 2023). More advanced and specific collusion strategies include alternating attacks and stealthy collusion. In alternating (on-off) attacks, malicious clients alternate between benign and malicious behavior to build reputation or evade history-based detection (Lewis et al., 2023). In stealthy collusion attacks, attackers coordinate to make their cumulative malicious impact significant while keeping individual updates close to benign ones to evade detection (Lyu et al., 2025). Such attacks aim for sparsity and stealthiness.
@@
3 Problem Formulation and Preliminaries
We consider a centralized FL setting in which, during each communication round, a central server broadcasts the current model weights to all clients. Each client then performs stochastic gradient descent on the loss function on its local dataset to compute an update. These local updates are sent back to the server, which aggregates them—most commonly through simple averaging—and applies a global gradient descent step scaled by a predefined learning rate. To enable a more precise formulation and analysis of the attacks, we assume that the server aggregates updates from all clients in every round—an assumption that commonly holds in cross-silo FL settings (Huang et al., 2024). We defer to future work the exploration of FL variants that either allow partial client participation or permit clients to perform several local updates before aggregation.
@@ 3 Problem Formulation and Preliminaries
with initialization $\hat{\theta}^{\mathcal{P}}_{0}=\theta_{0}$. Next, we introduce scenarios involving dictator clients in FL, including both single-dictator and multi-dictator cases. We describe how these clients modify their local updates to achieve their objectives. Specifically, a single dictator client aims to steer the global model’s updates and convergence to follow Eq. 2, while a group of coordinated dictator clients (collaborative dictators) seeks to enforce convergence toward Eq. 3, effectively overriding the standard FL update rule in Eq. 1.
4 Dictator Client Scenarios
In this section, we propose algorithms that enable clients to become dictators—retaining their own contributions to the global model while eliminating those of others. We begin with the case of a single dictator client in Section 4.1 and then extend to scenarios involving multiple collaborating dictator clients in Section 4.2. Figure 1 illustrates different dictator client scenarios compared with standard FL.
4.1 Single Dictator Client
In this section, we demonstrate how a single dictator client can craft its updates to entirely nullify the contributions of all other clients while preserving its own influence on the global model. We assume that the dictator client knows only the server’s learning rate and requires no additional information. Notably, as shown in Appendix C, even this assumption can be relaxed, as the learning rate can be numerically estimated after a single iteration. Suppose client $m\in\mathcal{N}$ such that $1\leq m\leq N$ is the designated dictator client and only knows server’s learning rate $\eta$. At iteration 0, the server broadcasts the initial model $\theta_{0}$ to all clients, which each use to compute their local gradients. Upon receiving these gradients, the server updates the global model as $\theta_{1}=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}$. In the next iteration, the server broadcasts $\theta_{1}$ to all clients. Each client except client $m$, computes and sends their gradient with respect to $\theta_{1}$. Meanwhile, client $m$ retains a local copy of the initial server model $\theta_{0}$ from the previous iteration. Using this, it computes a hypothetical model update, denoted by $\hat{\theta}^{m}_{1}$, which represents the model that would have resulted if only client $m$’s gradient had been used in the first iteration. This is computed as:
@@ 4.1 Single Dictator Client
The dictator client $m$ sends a carefully crafted update $M_{1}$ instead of its actual gradient $\nabla{\mathcal{L}_{m}(\theta_{1})}$ to delete the contribution of all other clients from the previous iteration and preserve only its own contribution. This manipulated update is defined as $M_{1}=\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})-\left(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{m}(\theta_{0})\right)}$. Here, the term $\frac{\theta_{0}-\theta_{1}}{\eta}$ reconstructs the aggregate gradient used by the server in the first round, allowing client $m$ to effectively cancel out the influence of all other clients while steering the update toward its own objective. We now analyze the updated global model $\theta_{2}$ after the server aggregates all client updates in the second iteration:
@@
$=\theta_{1}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})}-(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla{\mathcal{L}_{m}(\theta_{0}))}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{1})}})$
$=\theta_{0}-\eta\nabla{\mathcal{L}_{m}(\theta_{0})}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{1})}}).$
Now, substituting from Eq. 4, we can express $\theta_{2}$ as:
@@
This demonstrates that by sending the carefully crafted update $M_{1}$, client $m$ effectively nullifies the contributions of all other clients from the previous iteration while retaining its own gradient contribution. In doing so, the server’s model state is steered toward the single-client trajectory $\hat{\theta}^{m}_{1}$ instead of the standard FL update. To generalize this strategy for any round $t$, client $m$ maintains a local model $\hat{\theta}^{m}_{t}$, which is updated independently as:
$\hat{\theta}^{m}_{t}=\hat{\theta}^{m}_{t-1}-\eta\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1})}.$ (5)
We define $M_{t}$ as the update that the dictator client $m$ sends to the server at iteration $t$:
$M_{t}=\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}-\left(\frac{\hat{\theta}^{m}_{t-1}-\theta_{t}}{\eta}-\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1}})\right).$ (6)
@@
$=\theta_{t}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}-(\frac{\hat{\theta}^{m}_{t-1}-\theta_{t}}{\eta}-\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1}))}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}})$
$=\hat{\theta}^{m}_{t-1}-\eta\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1})}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}).$
Using Eq. 5, it follows that $\theta_{t+1}=\hat{\theta}^{m}_{t}-\eta\left(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}\right)$.
After $T$ rounds of training, the final model weights $\theta^{*}$ will be:
@@
This final expression shows that the dictator client drives the model toward its own trajectory $\hat{\theta}^{m}_{T+1}$, effectively overriding the influence of other clients up to a residual term. As shown in Eq. 8, the exact final weights under our method are given by $\theta^{*}=\hat{\theta}^{m}_{T+1}-\eta{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$, where $\hat{\theta}^{m}_{T+1}$ represents the weights for the final iteration if only the dictator client $m$ had participated in training, as if no other client existed. The residual term $\eta{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$ captures the contributions from the other clients in the final iteration. In practice, this term is negligible, as it stems from a single round of updates and has minimal impact on the final model—especially when the model produced by the dictator client is robust to such perturbations. Our empirical results in Section 6 further support the insignificance of this residual term on the dictator client’s objective. Algorithm 1 outlines the complete procedure for a client to act as a dictator.
4.2 Collaborative Dictator Clients

### v2

@@ Abstract
Federated learning (FL) [16] is a distributed learning paradigm in which model training is performed collaboratively by a set of clients. In centralized FL, a global server broadcasts the current model to all clients, each of which updates the model using its local dataset and sends back the resulting gradients to the server. The server then aggregates these gradients to update the global model. This approach accelerates training by distributing computation across multiple machines, while also enhancing data privacy since clients share only gradients, not their raw data. FL is especially well-suited for privacy-sensitive applications, such as training on confidential medical records across hospitals.
Despite its advantages, FL remains vulnerable to malicious behavior by the participating clients. Byzantine clients are adversarial participants that disrupt the training process by sending arbitrary or manipulated updates to the central server [8][3]. The presence of such adversaries can significantly degrade model performance, making Byzantine robustness a critical area of study [11, 24, 22, 5, 27, 28]. Moreover, several studies have demonstrated the possibility of backdoor attacks in FL via collusion attacks where multiple malicious clients coordinate their actions to inject hidden triggers into the global model in FL [13, 19, 25, 1]. These clients may exchange information and strategically craft updates that steer the aggregated model toward a compromised state.
@@
To demonstrate the feasibility of this threat, we develop a series of algorithms that enable malicious clients to achieve their goals within the defined constraints. Our theoretical findings are further supported by empirical results, which validate the effectiveness of the proposed attack strategies. Beyond isolated attacks, we also investigate complex and previously underexamined dynamics that arise among malicious clients themselves. For example, we examine scenarios in which all participants in the system act as dictators, as well as cases where collaborative dictator clients can betray one another within their own partnership. These scenarios reveal internal conflicts among adversaries and broaden the understanding of multi-agent adversarial behavior in FL. The practical implications of dictator clients are also discussed in more detail in Appendix E.
I Related Work
The distributed nature of FL, combined with the server’s limited visibility into local training processes, makes it vulnerable to various security threats posed by malicious or compromised clients [31]. In this section, we review existing literature across three major category of attacks—Byzantine attacks, backdoor attacks, and collusion attacks.
@@ I Related Work
Byzantine attacks pose a fundamental threat in distributed systems including FL, where a subset of clients, known as Byzantine clients, arbitrarily deviate from the prescribed protocol by submitting malicious or anomalous updates to the central server [8]. The goals of such attacks typically include degrading the global model’s performance or preventing convergence [3]. Attack strategies vary in complexity, ranging from simple approaches such as random noise injection or submitting zero gradients to more sophisticated methods like sign-flipping [20, 23]. Advanced attacks are often crafted to evade specific defenses, making them challenging to detect and mitigate [22, 2].
@@
Backdoor attacks (also known as Trojan attacks) are a more insidious threat in FL where attackers aim to embed hidden malicious behavior into the global model [4, 12]. An attacker, typically controlling one or more clients, manipulates their local dataset or model updates to create a ”backdoor trigger”—a specific pattern or feature (e.g., a small patch in an image, a specific phrase in text). The compromised global model performs normally on clean inputs but exhibits attacker-chosen behavior, such as misclassification, when the trigger is present. These attacks can be implemented through various strategies, including data poisoning, where labels are manipulated for samples containing the trigger, and model poisoning, where malicious updates are directly crafted to influence model behavior [1, 26]. Triggers may be static and predefined [1] or dynamically generated using optimization techniques to make them more subtle and difficult to detect [29]. Comprehensive surveys on backdoor attacks and defenses in FL can be found in [17].
@@
Collusion attacks occur when multiple malicious clients coordinate their actions to enhance the effectiveness of the attacks or bypass defenses designed for independent attackers. Colluding attackers can amplify the impact of Byzantine or backdoor attacks. For example, multiple Byzantine clients might coordinate their updates to overwhelm Byzantine-resilient aggregation rules that assume the number of attackers are limited [26]. Similarly, colluding clients can implement distributed backdoor attacks, where each attacker contributes a part of the malicious update, making individual contributions appear benign while collectively embedding a backdoor into the global model [15]. More advanced and specific collusion strategies include alternating attacks and stealthy collusion. In alternating (on-off) attacks, malicious clients alternate between benign and malicious behavior to build reputation or evade history-based detection [10]. In stealthy collusion attacks, attackers coordinate to make their cumulative malicious impact significant while keeping individual updates close to benign ones to evade detection [14]. Such attacks aim for sparsity and stealthiness.
@@
II Problem Formulation and Preliminaries
We consider a centralized FL setting in which, during each communication round, a central server broadcasts the current model weights to all clients. Each client then performs stochastic gradient descent on the loss function on its local dataset to compute an update. These local updates are sent back to the server, which aggregates them—most commonly through simple averaging—and applies a global gradient descent step scaled by a predefined learning rate. To enable a more precise formulation and analysis of the attacks, we assume that the server aggregates updates from all clients in every round—an assumption that commonly holds in cross-silo FL settings [6]. We defer to future work the exploration of FL variants that either allow partial client participation or permit clients to perform several local updates before aggregation.
@@ II Problem Formulation and Preliminaries
with initialization $\hat{\theta}^{\mathcal{P}}_{0}=\theta_{0}$. Next, we introduce scenarios involving dictator clients in FL, including both single-dictator and multi-dictator cases. We describe how these clients modify their local updates to achieve their objectives. Specifically, a single dictator client aims to steer the global model’s updates and convergence to follow Eq. (2), while a group of coordinated dictator clients (collaborative dictators) seeks to enforce convergence toward Eq. (3), effectively overriding the standard FL update rule in Eq. (1).
III Dictator Client Scenarios
In this section, we propose algorithms that enable clients to become dictators—retaining their own contributions to the global model while eliminating those of others. We begin with the case of a single dictator client in Section III-A and then extend to scenarios involving multiple collaborating dictator clients in Section III-B. Figure 3 (in Appendix A) illustrates different dictator client scenarios compared with standard FL.
III-A Single Dictator Client
In this section, we demonstrate how a single dictator client can craft its updates to entirely nullify the contributions of all other clients while preserving its own influence on the global model. We assume that the dictator client knows only the server’s learning rate and requires no additional information. Notably, as shown in Appendix H, even this assumption can be relaxed, as the learning rate can be numerically estimated after a single iteration. Suppose client $m\in\mathcal{N}$ such that $1\leq m\leq N$ is the designated dictator client and only knows server’s learning rate $\eta$. At iteration 0, the server broadcasts the initial model $\theta_{0}$ to all clients, which each use to compute their local gradients. Upon receiving these gradients, the server updates the global model as $\theta_{1}=\theta_{0}-\eta\sum\nolimits_{n=1}^{N}\nabla{\mathcal{L}_{n}(\theta_{0})}$. In the next iteration, the server broadcasts $\theta_{1}$ to all clients. Each client except client $m$, computes and sends their gradient with respect to $\theta_{1}$. Meanwhile, client $m$ retains a local copy of the initial server model $\theta_{0}$ from the previous iteration. Using this, it computes a hypothetical model update, denoted by $\hat{\theta}^{m}_{1}$, which represents the model that would have resulted if only client $m$’s gradient had been used in the first iteration. This is computed as:
@@ III-A Single Dictator Client
The dictator client $m$ sends a carefully crafted update $M_{1}$ instead of its actual gradient $\nabla{\mathcal{L}_{m}(\theta_{1})}$ to delete the contribution of all other clients from the previous iteration and preserve only its own contribution. This manipulated update is defined as:
$M_{1}=\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})-\left(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla\mathcal{L}_{m}(\theta_{0})\right)}.$
Here, the term $\frac{\theta_{0}-\theta_{1}}{\eta}$ reconstructs the aggregate gradient used by the server in the first round, allowing client $m$ to effectively cancel out the influence of all other clients while steering the update toward its own objective. We now analyze the updated global model $\theta_{2}$ after the server aggregates all client updates in the second iteration:
@@
$=\theta_{1}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})}-(\frac{\theta_{0}-\theta_{1}}{\eta}-\nabla{\mathcal{L}_{m}(\theta_{0}))}$
$\quad+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{1})}})$
$=\theta_{0}-\eta\nabla{\mathcal{L}_{m}(\theta_{0})}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{1})}$
$\quad+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{1})}}).$
Now, substituting from Eq. (4), we can express $\theta_{2}$ as:
@@
This demonstrates that by sending the carefully crafted update $M_{1}$, client $m$ effectively nullifies the contributions of all other clients from the previous iteration while retaining its own gradient contribution. In doing so, the server’s model state is steered toward the single-client trajectory $\hat{\theta}^{m}_{1}$ instead of the standard FL update. To generalize this strategy for any round $t$, client $m$ maintains a local model $\hat{\theta}^{m}_{t}$, which is updated independently as $\hat{\theta}^{m}_{t}=\hat{\theta}^{m}_{t-1}-\eta\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1})}$. We define $M_{t}$ as the update that the dictator client $m$ sends to the server at iteration $t$ as:
$M_{t}=\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}-\left(\frac{\hat{\theta}^{m}_{t-1}-\theta_{t}}{\eta}-\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1}})\right).$
@@
$=\theta_{t}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}-(\frac{\hat{\theta}^{m}_{t-1}-\theta_{t}}{\eta}-\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1}))}$
$\quad+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}})$
$=\hat{\theta}^{m}_{t-1}-\eta\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t-1})}-\eta(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}$
$\quad+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}).$
Substituting $\hat{\theta}^{m}_{t}$, it follows that $\theta_{t+1}=\hat{\theta}^{m}_{t}-\eta\left(\nabla{\mathcal{L}_{m}(\hat{\theta}^{m}_{t})}+{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}\right)$. After $T$ rounds of training, the final model weights $\theta^{*}$ will be:
@@
This final expression shows that the dictator client drives the model toward its own trajectory $\hat{\theta}^{m}_{T+1}$, effectively overriding the influence of other clients up to a residual term. As shown in Eq. (6), the exact final weights under our method are given by $\theta^{*}=\hat{\theta}^{m}_{T+1}-\eta{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$, where $\hat{\theta}^{m}_{T+1}$ represents the weights for the final iteration if only the dictator client $m$ had participated in training, as if no other client existed. The residual term $\eta{\sum\nolimits_{n=1,n\neq{m}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$ captures the contributions from the other clients in the final iteration. In practice, this term is negligible, as it stems from a single round of updates and has minimal impact on the final model—especially when the model produced by the dictator client is robust to such perturbations. Our empirical results in Section IV further support the insignificance of this residual term on the dictator client’s objective. Algorithm 1 (in Appendix B) outlines the complete procedure for a client to act as a dictator.
Method MNIST CIFAR-10
[0,1] [2,3] [4,5] [6,7] [8,9] [0,1] [2,3] [4,5] [6,7] [8,9]
Regular FL 96.18 79.25 66.84 88.12 66.38 39.04 12.51 31.07 23.74 52.59
Dictator client: 11 99.63 0.00 0.00 0.00 0.00 73.65 0.00 0.00 0.00 0.00
Dictator client: 22 0.00 93.92 0.00 0.00 0.00 0.00 65.19 0.00 0.00 0.00
Dictator client: 33 0.00 0.00 97.43 0.00 0.00 0.00 0.00 66.51 0.00 0.00
Dictator client: 44 0.00 0.00 0.00 98.91 0.00 0.00 0.00 0.00 73.98 0.00
Dictator client: 55 0.00 0.00 0.00 0.00 94.42 0.00 0.00 0.00 0.00 77.06
TABLE I: Performance of the global model on each local dataset for MNIST and CIFAR-10 datasets and the single dictator client and collaborative dictator clients scenarios.
III-B Collaborative Dictator Clients